OpenStudy (theedwardsfamily):

Help!!! Will medal! For which operations is the set {–1, 0, 1} closed? Choose all answers that are correct. A. addition B. division C. multiplication D. subtraction

OpenStudy (zzr0ck3r):

do you know what closed means?

OpenStudy (zzr0ck3r):

1+1=2 is not in the set 1/0 is not defined -1-1 = -2 is no in the set

OpenStudy (zzr0ck3r):

@TheEdwardsFamily

OpenStudy (theedwardsfamily):

no I dont know

OpenStudy (zzr0ck3r):

so a set is closed with respect to an operation if for ALL elements a,b in the set we have that (a operator b) is in the set

OpenStudy (zzr0ck3r):

so can you take any two elements of the set and add them together and get another element in the set?

OpenStudy (theedwardsfamily):

So a and b are correct?

OpenStudy (theedwardsfamily):

umm con fused

OpenStudy (theedwardsfamily):

brb

OpenStudy (zzr0ck3r):

? lol

OpenStudy (zzr0ck3r):

so you would agree that 1 is in the set right? then IF it is closed with respect to addition we have 1+1 is in the set, is that true?

OpenStudy (theedwardsfamily):

All im thinking is numbers whirling in my head! Its annoying!:(

OpenStudy (theedwardsfamily):

@KamiBug are you good at math??

OpenStudy (theedwardsfamily):

@Nnesha I need you!

OpenStudy (theedwardsfamily):

@Directrix

OpenStudy (theedwardsfamily):

@KamiBug

Nnesha (nnesha):

okay so take one number like 1 first add this one into this set values {-1,0,1} like 1+1 = 2 but 2 is not in the set then subtract -1-1=?? -2 is not there

Nnesha (nnesha):

zzr0ck3r already explained everything :)

OpenStudy (theedwardsfamily):

but i dont understand

OpenStudy (zzr0ck3r):

ok do you agree that 1 is in the set?

OpenStudy (theedwardsfamily):

i guess

OpenStudy (zzr0ck3r):

the set is \(\{-1,0,1\}\)

OpenStudy (zzr0ck3r):

is 1 in it?

OpenStudy (zzr0ck3r):

I don't want you to guess...

OpenStudy (zzr0ck3r):

do you see a 1?

OpenStudy (theedwardsfamily):

yea

OpenStudy (zzr0ck3r):

ok we are talking about part a) addition is 1+1 = 2 in the set?

OpenStudy (zzr0ck3r):

I am asking if 2 is in the set.

OpenStudy (theedwardsfamily):

no

OpenStudy (zzr0ck3r):

ok so since we did 1+1=2 and 2 is not in the set, we have that it is not closed with respect to +

OpenStudy (theedwardsfamily):

ok

OpenStudy (zzr0ck3r):

you could have used -1 also -1 is in the set, but -1 + -1 = -2 is not in the set

OpenStudy (theedwardsfamily):

uh-huh

OpenStudy (zzr0ck3r):

ok so addition is not closed what about subtraction?

OpenStudy (theedwardsfamily):

yes?

OpenStudy (zzr0ck3r):

why?

OpenStudy (theedwardsfamily):

guessed

OpenStudy (zzr0ck3r):

ok im done. good luck man.

Nnesha (nnesha):

okay listen

OpenStudy (theedwardsfamily):

Not a man

OpenStudy (zzr0ck3r):

what ever you are, I am done lol @Nnesha good luck

Nnesha (nnesha):

take a number like oneafter subtraction you get -2 is that on the set